# An Optimism-based Approach to Online Evaluation of Generative Models

### v1

@@ A.4 Concentration of Mean Vector and Covariance Matrix
A.5 Dimension-free Concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
Theorem A.3 (Dimension-free concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$).
Under the same conditions in Theorem A.1, we have that
@@
Let $\eta=\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}$. We have that
@@
$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\widehat{\Sigma}_{g}^{n}+\eta I)B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
@@
where the first inequality holds by Lemma E.10, the second inequality holds by Lemma E.11 and the fact that both matrices $B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B\succeq\bm{0},B^{\top}\widehat{\Sigma}_{g}^{n}B\succeq\bm{0}$ are PSD, and the last inequality holds by Lemma E.10 and $B^{\top}(2\eta I)B\succeq B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B$. By the same analysis, we can also derive that
@@
Appendix B Proofs in Section 6: IS Evaluation
B.1 Proof of Theorem 6.1: Generator-Dependent Optimistic IS
@@ Appendix B Proofs in Section 6: IS Evaluation
By Lemma B.2, we have that with probability at least $1-\frac{\delta}{2}$, ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$. To show Theorem 6.1, it suffices to show that
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}-\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)$
with probability at least $1-\frac{\delta}{2}$, which holds by Theorem E.4. Therefore, we conclude the proof. ∎
B.2 Proof of Theorem 6.2: Regret of IS-UCB
@@
By Theorem 6.1 and union bound over $T$ steps, with probability at least $1-\delta$, it holds that $\widehat{\textup{IS}}_{g}\geq\textup{IS}_{g}$ for all steps $t\in[T]$ and $g\in[G]$, where $\widehat{\textup{IS}}_{g}$ is given by Equation (14). For convenience, we denote by $\widehat{\textup{IS}}_{g_{t}}$ the optimistic IS of generator $g_{t}$ computed at the $t$-th step. Hence, we have that
@@
where
${\mathcal{B}}_{g_{t}}:=\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right).$
Let ${\mathcal{T}}_{1},{\mathcal{T}}_{2}\subset[T]$ be two disjoint set of steps such that ${\mathcal{T}}_{2}=[T]\backslash{\mathcal{T}}_{1}$. Specifically, ${\mathcal{T}}_{1}$ contains steps where no element of the empirical marginal class distribution $\widetilde{p}_{Y_{g_{t}}}$ is clipped to $e^{-1}$, i.e., $|e^{-1}-\widetilde{p}_{Y_{g_{t}}}[j]|\geq\bm{\epsilon}_{g_{t}}[j]$ for all $[j]\in d$. Let $\textup{IS}^{*}=\operatorname*{arg\,max}_{g\in[G]}\textup{IS}_{g}$ denote the optimal Inception score. Hence, we have that
@@
$\leq$ $\widetilde{O}\left(e^{C}\cdot\sum_{t=1}^{T}\left(\max\left\{dm,\frac{d\log d}{2}\right\}\sqrt{\frac{1}{N_{g_{t}}}}+\frac{C^{\prime}}{N_{g_{t}}}\right)\right),$ (31)
where $m=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, and $C=O(de^{-1}+C^{\prime})$ is the upper bound of ${\mathcal{E}}(\widehat{p}_{Y_{g_{t}}})-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}$. For the second part, note that the number of steps where at least one element of $\widetilde{p}_{Y_{g_{t}}}$ is clipped to at most $\widetilde{O}(G\Delta^{-2})$, where $\Delta=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$. Therefore, we have that
$\textup{Regret}(T)\leq\widetilde{O}\left(dG\Delta^{-2}\log d+e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}\right),$
@@
B.3 Optimistic marginal class distribution
Theorem B.1 (Optimistic marginal class distribution).

### v3

@@ A.4 Concentration of Mean Vector and Covariance Matrix
A.5 Concentration Bound for Norm-Bounded Embeddings
We extend the concentration bound for norm-bounded embeddings.
Theorem 5.
Assume for any generator $g$, the (random) embedding $f(X_{g})$ is $L2$-norm bounded, i.e., $\|f(X_{g})\|_{2}\leq C$ for some positive number $C>0$, and the covariance matrix $\Sigma_{\textup{r}}$ of the real data is positive definite. Then, with probability at least $1-\delta$, we have
$\widetilde{\textup{FD}}_{g}^{n}-{\mathcal{C}}^{n}_{g}\leq\textup{FD}_{g},$
where
${\mathcal{C}}^{n}_{g}:=2(\Delta_{1}^{n}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\Delta_{1}^{n}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\Delta_{2}^{n}}+4C^{2}\sqrt{\frac{1}{2n}\log\left(\frac{3}{\delta}\right)},$
and $\Delta_{1}^{n}$ and $\Delta_{2}^{n}$ are defined in (28) and (29), respectively.
Proof.
The proof is similar to Theorem 1. Recall that for any generator $g$, we have that
$\textup{FD}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
$\widetilde{\textup{FD}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  We derive
$\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}\right|\leq(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+2\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
(2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.  By Lemma 2, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
$\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}\left[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$
(3) Bound $\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]$.  Note that
$\left|\textup{Tr}[\Sigma_{g}]-\textup{Tr}[\widehat{\Sigma}_{g}^{n}]\right|\leq\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}.$
As $\|f(X_{g})\|_{2}\leq C$, we have $\|f(X_{g})-\mu_{g}\|_{2}^{2}\leq(\|f(X_{g})\|_{2}+\|\mu_{g}\|_{2})^{2}\leq(2C)^{2}$. By Hoeffding’s inequality, we have that with probability at least $1-\frac{\delta}{3}$, it holds that
$\left|\frac{1}{n}\sum_{i=1}^{n}\|f(X^{i})-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]\right|\leq 4C^{2}\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}$
Therefore, we have
$\left|\textup{FD}_{g}-\widetilde{\textup{FD}}_{g}^{n}\right|\leq 2(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}+4C^{2}\sqrt{\frac{1}{2n}\log\left(\frac{3}{\delta}\right)}.$ (27)
Next, we analyze the concentration error of the $L2$-norm error. Since
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}=\underbrace{\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{ concentration}}+\underbrace{\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{expected $L$2-norm error}},$
the expected $L$2-norm error is given by
$\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]=\frac{\textup{Tr}[\Sigma_{g}]}{n}.$
Additionally, we have that $|\widehat{\mu}_{g}^{n}[i]-\mu_{g}[i]|\leq 2C$ for the $i$-th entry where $i\in[d]$. Hence, by Hoeffding’s inequality and a union bound over all dimensions, with probability at least $1-\frac{\delta}{3}$, it holds that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]\leq\frac{2dC^{2}}{n}\log\left(\frac{6d}{\delta}\right).$
Hence, we have
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{2dC^{2}}{n}\log\left(\frac{6d}{\delta}\right)+\frac{\textup{Tr}[\Sigma_{g}]}{n}}:=\Delta_{1}^{n}$ (28)
Further, we have that with probability at least $1-\frac{\delta}{3}$, it holds that
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+(\Delta_{1}^{n})^{2}:=\Delta_{2}^{n}$ (29)
for $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, which concludes the proof.
A.6 Auxiliary Definitions and Lemmas
Lemma 2 (Error decomposition for $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$).
Under the same conditions in Theorem 4, we have that
@@ A.6 Auxiliary Definitions and Lemmas
Let $\eta:=\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}$ and $\Sigma_{\textup{r}}:=BB^{\top}$ denote the Cholesky decomposition of $\Sigma_{\textup{r}}$, where $B$ is invertible by the assumption $\Sigma_{\textup{r}}\succ\bm{0}$. Note that matrix $A\Sigma_{\textup{r}}=ABB^{\top}$ has the same set of eigenvalues with $B^{\top}AB$. We have that
@@
$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\widehat{\Sigma}_{g}^{n}+\eta I)B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$ (Lemma 7)
@@
where the second inequality holds by Lemma 8 and the fact that $B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B,B^{\top}\widehat{\Sigma}_{g}^{n}B\succeq\bm{0}$ are PSD, and the last inequality holds by $B^{\top}(2\eta I)B\succeq B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B$. By the same analysis, we can also derive that
@@
Definition 1 (Sub-Gaussian random variable).
A random variable $X$ with mean $\mu=\mathbb{E}[X]$ is sub-Gaussian with (positive) parameter $\sigma$, which is denoted by $X\in\textup{SG}(\sigma)$, if
$\mathbb{E}[\exp(s(X-\mu))]\leq\exp\left(\frac{s^{2}\sigma^{2}}{2}\right)$
for all $s\in{\mathbb{R}}$.
Definition 2 (Sub-Gaussian random vector).
A random vector $X\in{\mathbb{R}}^{d}$ is a sub-Gaussian random vector with (positive) parameter $\sigma$ if $y^{\top}X\in\textup{SG}(\sigma)$. Particularly, the multivariate Gaussian $X\sim{\mathcal{N}}(0,\Sigma)\in\textup{SG}(\sqrt{\|\Sigma\|_{2}})$.
Definition 3 (Sub-exponential random variable).
A random variable $X$ with mean $\mu=\mathbb{E}[X]$ is sub-exponential with (non-negative) parameters $(\nu,\alpha)$, which denoted by $X\in\textup{SE}(\nu,\alpha)$, if
$\mathbb{E}[\exp(s(X-\mu))]\leq\exp\left(\frac{s^{2}\nu^{2}}{2}\right)$
for all $|s|<\frac{1}{\alpha}$. Moreover, we have
${\mathbb{P}}[X-\mathbb{E}[X]\geq t]\leq\begin{cases}\exp(-\frac{t^{2}}{2\nu^{2}})&\textup{if }0\leq t\leq\frac{\nu^{2}}{\alpha},\\ \exp(-\frac{t}{2\alpha})&\textup{for }t>\frac{\nu^{2}}{\alpha}.\end{cases}$
Equivalently, with probability at least $1-\delta$, it holds that
$|X-\mathbb{E}[X]|\leq\max\left\{\sqrt{2\nu^{2}\log\left(\frac{2}{\delta}\right)},2\alpha\log\left(\frac{2}{\delta}\right)\right\}.$
Lemma 3 (Weighted sum of sub-exponential r.v.’s).
Let $X_{1},\cdots,X_{n}$ be $n$ independent random variables, where $X_{i}\in\textup{SE}(\nu_{i},\alpha_{i})$ is sub-exponential with mean $\mathbb{E}[X_{i}]=\mu_{i}$ and parameters $(\nu_{i},\alpha_{i})$. Then, for any $W=(w_{1},\cdots,w_{n})^{\top}\in{\mathbb{R}}^{n}_{+}$, the (non-negative) weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$ is sub-exponential with parameters $(\sqrt{\sum_{i=1}^{n}w_{i}^{2}\nu_{i}^{2}},\max_{i}\{w_{i}\cdot\alpha_{i}\})$.
@@
By Definition 3, we have
$\mathbb{E}[\exp(s\cdot w_{i}(X_{i}-\mu_{i}))]\leq\exp\left(\frac{s^{2}w_{i}^{2}\nu_{i}^{2}}{2}\right)$
for any $i\in[n]$ and $|s|\cdot w_{i}<\frac{1}{\alpha_{i}}$ Further, since $X_{1},\cdots,X_{n}$ are independent, it holds that
$\mathbb{E}[\exp(s\cdot\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{i}))]\leq\prod_{i=1}^{n}\exp\left(\frac{s^{2}w_{i}^{2}\nu_{i}^{2}}{2}\right)=\exp\left(\frac{s^{2}\left(\sqrt{\sum_{i=1}^{n}w_{i}^{2}\nu_{i}^{2}}\right)^{2}}{2}\right)$
for any $|s|<\min_{i}\frac{1}{w_{i}\cdot\alpha_{i}}$, which concludes the proof. ∎
Lemma 4 (Squared L2-norm of multivariate Gaussian).
Let $X\sim{\mathcal{N}}(\bm{0},\Sigma)$ be any $d$-dimensional multivariate Gaussian with zero mean. Then, the r.v. $\|X\|_{2}^{2}\sim\sum_{i=1}^{d}\lambda_{i}\chi^{2}_{1}$ is a weighted sum of i.i.d. chi-squared random variables with one degree of freedom, where $\{\lambda_{i}\}_{i=1}^{d}$ are the eigenvalues of the covariance matrix $\Sigma$. Further, let $X_{1},\cdots,X_{n}\sim{\mathcal{N}}(\bm{0},\Sigma)$ denote i.i.d. samples. Then, with probability at least $1-\delta$, it holds that
$\left|\frac{1}{n}\sum_{i=1}^{n}\|X_{i}\|_{2}^{2}-\mathbb{E}[\|X\|_{2}^{2}]\right|\leq\max\left\{\sqrt{\frac{8\cdot\textup{Tr}[\Sigma^{2}]}{n}\log\left(\frac{2}{\delta}\right)},\frac{8\|\Sigma\|_{2}}{n}\log\left(\frac{2}{\delta}\right)\right\},$ (31)
where $\mathbb{E}[\|X\|_{2}^{2}]:=\sum_{i=1}^{d}\lambda_{i}=\textup{Tr}[\Sigma]$.
@@
Let $\Sigma=Q\Lambda Q^{\top}$ be the eigendecomposition of the covariance matrix, where we define $\Lambda:=\textup{diag}(\lambda_{1},\cdots,\lambda_{d})$. Hence, we have $X\sim QZ$, where $Z\sim{\mathcal{N}}(\bm{0},\Lambda)$ is isotropic multivariate Gaussian. Hence, we derive
$\|X\|_{2}^{2}=(QZ)^{\top}QZ=Z^{\top}Q^{\top}QZ=Z^{\top}Z=\sum_{i=1}^{d}\lambda_{i}\chi_{1}^{2}.$
Since $\chi_{1}^{2}\in\textup{SE}(2,4)$ is sub-exponential with parameters $(2,4)$, by Lemma 3 (with $w_{i}=\lambda_{i}$), we have $\|X\|_{2}^{2}\in\textup{SE}(2\sqrt{\textup{Tr}[\Sigma^{2}]},4\|\Sigma\|_{2})$ and has mean $\mathbb{E}[\|X\|_{2}^{2}]=\textup{Tr}[\Sigma]$. We conclude the proof by Lemma 5. ∎
Lemma 5 (Bernstein-type bound for sub-exponential r.v.).
Let $X_{1},\cdots,X_{n}$ be $n$ i.i.d. random variables, where each $X_{i}\in\textup{SE}(\nu,\alpha)$. Then, with probability at least $1-\delta$, it holds that
$\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mathbb{E}[X]\right|\leq\max\left\{\sqrt{\frac{2\nu^{2}}{n}\log\left(\frac{2}{\delta}\right)},\frac{2\alpha}{n}\log\left(\frac{2}{\delta}\right)\right\}$
Proof.
By Lemma 3, we have $\frac{1}{n}\sum_{i=1}^{n}X_{i}\in\textup{SE}(\frac{\nu}{\sqrt{n}},\frac{\alpha}{n})$, which concludes the proof. ∎
Lemma 6 (Dimension-free concentration of sample covariance matrix (Lounici,, 2014; Koltchinskii and Lounici,, 2017; Zhivotovskiy,, 2022)).
Let $X_{1},\cdots,X_{n}$ are i.i.d sub-Gaussian random vectors with zero mean and covariance matrix $\Sigma$. Let
$\bm{r}(\Sigma):=\frac{\textup{Tr}(\Sigma)}{\|\Sigma\|_{2}}$ (32)
$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}-\Sigma\right\|_{2}\leq 20\kappa^{2}\|\Sigma\|_{2}\left(\sqrt{\frac{4\bm{r}(\Sigma)+\log(1/\delta)}{n}}\right)$ (33)
for $n\geq 4\bm{r}(\Sigma)+\log(1/\delta)$. Here, the parameter $\kappa$ is a constant such that $\|y^{\top}X\|_{\psi_{2}}\leq\kappa\sqrt{y^{\top}\Sigma y}$, where $\|Z\|_{\psi_{2}}:=\inf\{t>0:\mathbb{E}[\exp(Z^{2}/t^{2})]\leq 2\}$ is the sub-Gaussian norm.
Lemma 7.
Proof.
Lemma 8.
Proof.
$\leq$ $\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}],$
where the last inequality holds by $A^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ and $B^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ (see Lemma 9). For general PSD matrices $A$ and $B$, note that
Lemma 9.
Proof.
Appendix B PROOFS IN SECTION LABEL:SEC:6: IS-BASED EVALUATION
B.1 Proof of Theorem 2: Generator-Dependent Optimistic IS
Proof.
By Lemma 10, we have that with probability at least $1-\frac{\delta}{2}$, ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$. To show Theorem 2, it suffices to show that
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4d}{\delta}\right)}-\frac{7\log d}{3(n-1)}\log\left(\frac{4d}{\delta}\right)$
with probability at least $1-\frac{\delta}{2}$, which holds by Theorem 12 and the fact that $H_{g}(Y_{g}|X_{g})\leq\log d$. Therefore, we conclude the proof. ∎
B.2 Regret of IS-UCB
Theorem 6 (Regret of IS-UCB).
$\textup{Regret}(T)\leq\widetilde{O}\left(G\Delta^{-2}\log d+e^{C}\cdot\left((dm+\log d)\cdot\sqrt{T}+\log T\right)\right),$ (36)
where $C=\widetilde{O}(d)$, $m:=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, $\Delta:=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$, and $u(x)=-x\log x$.
Proof.
By Theorem 2 and union bound over $T$ steps, with probability at least $1-\delta$, it holds that $\widehat{\textup{IS}}_{g_{t}}\geq\textup{IS}_{g_{t}}$ for all steps $t\in[T]$ and $g\in[G]$, where $\widehat{\textup{IS}}_{g}$ is given by Equation (13). For convenience, we denote by $\widehat{\textup{IS}}_{g_{t}}$ the optimistic IS of generator $g_{t}$ computed at the $t$-th step. Hence, we have that
@@ B.2 Regret of IS-UCB
where we denote
${\mathcal{B}}_{g_{t}}:=\sqrt{\frac{2\widehat{V}_{g}^{n_{t}}(H(Y_{g}|X_{g}))}{n_{t}}\log\left(\frac{4Td}{\delta}\right)}+\frac{7\log d}{3(n_{t}-1)}\log\left(\frac{4Td}{\delta}\right)$
for convenience. Let ${\mathcal{T}}_{1},{\mathcal{T}}_{2}\subset[T]$ be two disjoint set of steps such that ${\mathcal{T}}_{2}=[T]\backslash{\mathcal{T}}_{1}$. Specifically, ${\mathcal{T}}_{1}$ contains steps where no element of the empirical marginal class distribution $\widetilde{p}_{Y_{g_{t}}}$ is clipped to $e^{-1}$, i.e., $|e^{-1}-\widetilde{p}_{Y_{g_{t}}}[j]|\geq\bm{\epsilon}_{g_{t}}[j]$ for all $[j]\in d$. Let $\textup{IS}^{*}=\operatorname*{arg\,max}_{g\in[G]}\textup{IS}_{g}$ denote the optimal Inception score. Hence, we have that
@@
$\leq$ $\widetilde{O}\left(e^{C}\cdot\sum_{t=1}^{T}\left((dm+\log d)\sqrt{\frac{1}{N_{g_{t}}}}+\frac{\log d}{N_{g_{t}}}\right)\right),$ (37)
where $m=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, and $C=\widetilde{O}(d)$ is the upper bound of ${\mathcal{E}}(\widehat{p}_{Y_{g_{t}}})-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}$. For the second part, note that the number of steps where at least one element of $\widetilde{p}_{Y_{g_{t}}}$ is clipped to at most $\widetilde{O}(G\Delta^{-2})$, where $\Delta=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$. Therefore, we have that
$\textup{Regret}(T)\leq\widetilde{O}\left(G\Delta^{-2}\log d+e^{C}\cdot\left((dm+\log d)\cdot\sqrt{T}+\log T\right)\right),$
@@
B.3 Data-dependent optimistic marginal class distribution
Lemma 10 (Optimistic marginal class distribution).
satisfies that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
Proof.
It suffices that show that $\bm{\epsilon}_{g}^{n}[j]\geq|\widehat{p}_{Y_{g}}^{n}[j]-p_{Y_{g}}[j]|$ for all $j\in[d]$ with probability at least $1-\frac{\delta}{2}$. We evoke Lemma 12 and conclude the proof. ∎
B.4 Auxiliary Definitions and Lemmas
Lemma 11 (Optimistic marginal class distribution).